can you please help me with this problem {3a+b+c=7
{ a+3b-c=13
{ b=2a-1

Answer by venugopalramana(3286)

3a+b+c=7..........(1)
a+3b-c=13........(2)
b=2a-1..........(3)...substituting value of b from eqn.(3) in eqns.(1)and(2)
we get...3a+2a-1+c=7...or...5a+c=7+1=8..............(4)
a+3*(2a-1)-c=13...or...a+6a-3-c=13...or...7a-c=13+3=16.........(5)
eqn.(4)+eqn.(5)...12a=8+16=24..or..a=24/12=2....substituting in (3)..we get
b=2*2-1=3...substituting in (4)we get..5*2+c=8..or..c=8-10=-2..
so a=2;b=3;c=-2.